OpenStudy (anonymous):

need help with 3 word problems please 1. The Divine Chocolate Candy store wants to make a new treat consisting of milk chocolate and hazelnuts. They want to make a 10 pound box that sells for $5 per pound. If hazelnuts sell for $6 per pound and chocolate sells for $4 per pound, how pounds of each must they mix to get their desired product? 2. A chemist wants to mix a 64% alcohol solution with 7 liters of 51% alcohol solution to produce a solution that is 62% alcohol. How many liters of the 64% alcohol solution should be used? Round to two decimal places. 3. Rick can mop the flo

OpenStudy (anonymous):

1. Solve the following equations for the pound weight of h and c, hazelnuts and chocolate.\[\{6h+4c=50,h+c=10\}\]\[\{h=5,c=5\} \] 2. Let x be the liter volume of 64% alcohol solution to be added. Solve the following for x:\[x \frac{64}{100}+7*\frac{51}{100}=(x+7)\frac{62}{100}\]\[x=\frac{77}{2}=38.5 \text{ liters} \]